# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v2

@@ B.2 Proof of Proposition 2
    Therefore, we have that $\mathcal{F}(\pi)=\lim_{n\to+\infty}\mathcal{F}_{\alpha_{n}}^{\eta_{n}}(\pi_{n})$, and $\{\mathcal{F}_{\alpha}^{\eta}:\,\alpha,\eta\geq 0\}$ is a $\Gamma$-system. In addition, for any $\alpha,\eta>0$, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}$ by Proposition 1. Using Proposition 8, we get that $\pi^{\star}$ is a minimizer of $\mathcal{F}$.
@@ Appendix C Subdifferential of $\mathcal{F}_{\alpha}^{\eta}$
$+\left\|\frac{(\lambda\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\nabla\varphi(x))(\lambda\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\nabla\varphi(x))^{T}}{(\mathcal{L}(\pi,x)+\eta)^{2}}\right\|$
@@
$\geq-\mathtt{C}_{0}\left\|t_{n}(x)-x\right\|^{2}.$
@@
$\;\phantom{\geq}-\mathtt{C}_{0}\left\|t_{n}(x)-x\right\|^{2}.$
Let us consider first
@@
$-\mathtt{C}_{0}\left\|t_{n}(x)-x\right\|^{2}\to 0.$
@@
$\lim_{n\to\infty}\{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\geq 0.$ (35)
In addition, since $\mathrm{k}$ is Lipschitz continuous with constant $\mathtt{M}$ we get that for any $\pi_{1},\pi_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{p}$ $|\mathcal{L}(\pi_{1},x)-\mathcal{L}(\pi_{2},x)|\leq\lambda\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}),$ and therefore
@@
By (35), this family satisfies $\liminf_{n\to\infty}f_{n}(x)\geq 0$. Consider the second derivative of (37)
@@
$\left\|\nabla h_{n}(x)\right\|$ $=\frac{\left\|\int_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(z)-\int\limits_{\mathbb{R}^{d}}\langle\nabla_{1}\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right\|}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq\frac{\int_{\mathbb{R}^{d}}\left\|\nabla_{1}\mathrm{k}(x,t_{n}(z))-\nabla_{1}\mathrm{k}(x,z)\right\|\mathrm{d}\pi(z)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
@@
$\leq(1+\lambda)\mathtt{M}\frac{\left(\int_{\mathbb{R}^{d}}\left\|t_{n}(z)-z\right\|^{2}\mathrm{d}\pi(z)\right)^{1/2}}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq\mathtt{M}(1+\lambda),$
@@
Hence, the sequence of functions is equicontinuous and we can apply Fatou’s Lemma for weakly converging measures [26, Theorem 4.1] to obtain
$\liminf_{n\to\infty}\frac{\int\limits_{\mathbb{R}^{d}}\!g^{\prime}(\mathcal{L}(\pi,x))\!\left(\!\mathcal{L}(t_{n\#}\pi,x)\!-\!\mathcal{L}(\pi,x)\!-\!\int\limits_{\mathbb{R}^{d}}\!\langle\xi_{2}(x,z),t_{n}(z)\!-\!z\rangle\mathrm{d}\pi(z)\!\right)\!\mathrm{d}t_{n\#}\pi(x)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}\geq 0,$
@@
$\partial_{\mathrm{s}}\mathrm{H}(\pi)=\{x\mapsto\nabla\log\pi(x)\}.$ (39)
@@ D.1 Proof of Proposition 3
To prove existence and uniqueness of the solution of (9) we use standard tools for McKean–Vlasov processes. In particular, we only need to show that the drift of (9), denoted by $b:\mathbb{R}^{d}\times\mathcal{P}_{1}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, given for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ by (13) is Lipschitz continuous. The remainder of the proof is classical and is omitted, see for instance [53, Theorem 1.1] or [16, Appendix B.2] for a recent result on a similar scheme.
@@ D.2 Proof of Proposition 4
Under Assumption 1, Proposition 1(a) ensures that the functional $\mathcal{F}^{\eta}$ is convex and lower bounded. In addition, Assumption 3 guarantees that $U$ is sufficiently regular to satisfy [32, Assumption 2.2]. To see this, we can use Assumption 3(c) and use the Cauchy–Schwarz inequality to obtain
@@
with $b^{\eta}$ in (10). As shown in (43), $D_{\nu}\mathcal{F}^{\eta}$ is Lipschitz continuous and thus $\mathcal{F}^{\eta}$ is continuously differentiable. In addition, using the fact that under Assumption 1 $\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ and Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]), we have that $D_{\nu}\mathcal{F}^{\eta}$ is $\mathrm{C}^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$ for all fixed $\nu\in\mathcal{P}(\mathbb{R}^{d})$. Finally, we need to show that
@@
$+\frac{\lambda}{\eta}\left\|x-x_{n}\right\|+\frac{\lambda^{2}\mathtt{M}^{2}}{\eta^{2}}\mathbf{W}_{1}(\nu_{n},\nu),$
@@
$\qquad\leq\frac{\lambda}{\eta}\left\|x-x_{n}\right\|+\left(\mathtt{M}_{1}+\frac{\lambda^{2}\mathtt{M}^{2}}{\eta^{2}}\right)\mathbf{W}_{1}(\nu_{n},\nu).$
@@
for some finite $C$. Since $\lim_{n\to+\infty}(x_{n},\nu_{n})=(x,\nu)$ implies $\left\|x-x_{n}\right\|\to 0$ and $\mathbf{W}_{1}(\nu_{n},\nu)\to 0$, we have that $\nabla b(x,\nu)$ is jointly continuous. Then, the result follows from [32, Theorem 2.11].
@@ E.1 Proof of Proposition 5
where we used the ordering of Wasserstein distances (see, e.g. [57, Remark 6.6]) for the first inequality and the definition of $\mathbf{W}_{2}$ for the second, we have that
@@
The existence and strong uniqueness of a solution to (11) is then a straightforward consequence of the above and [38, Chapter 5, Theorem 2.9 and 2.5].
@@
$\ \ \leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|)+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2})+\alpha\mathtt{L}\left\|x_{1}-x_{2}\right\|$
@@
where the last inequality follows using the ordering of Wasserstein distances (see, e.g. [57, Remark 6.6]) to obtain $W_{1}(\pi_{s},\pi_{s}^{N})\leq W_{2}(\pi_{s},\pi_{s}^{N})$ and the exchangeability of $\{(\bm{X}_{t}^{k,N})_{t\geq 0}\}_{k=1}^{N}$ to obtain
$\mathbb{E}\left[\mathbf{W}_{2}(\pi_{s},\pi_{s}^{N})\right]$ $\leq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\mathbf{W}_{2}(\pi_{s},\delta_{X_{s}^{i,N}})\right]$ (47)
$\leq\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\left\|\bm{X}_{s}-\bm{X}_{s}^{i,N}\right\|\right]\leq\mathbb{E}\left[\left\|\bm{X}_{s}-\bm{X}_{s}^{1,N}\right\|\right].$

### v3

@@ B.2 Proof of Proposition 2
    Therefore, we have that $\mathcal{F}(\pi)=\lim_{n\to+\infty}\mathcal{F}_{\alpha_{n}}^{\eta_{n}}(\pi_{n})$, and $\{\mathcal{F}_{\alpha}^{\eta}:\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})\to\mathbb{R}^{+},\ s.t.\ \alpha,\eta\geq 0\}$ is a $\Gamma$-system. In addition, for any $\alpha,\eta>0$, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}$ by Proposition 1, which under our assumptions, satisfies $\pi_{\alpha,\eta}^{\star}\in\mathcal{Q}_{2}^{ac}(\mathbb{R}^{d})$. Using Proposition 8, we get that $\pi^{\star}$ is a minimizer of $\mathcal{F}$.
@@ Appendix C Subdifferential of $\mathcal{F}_{\alpha}^{\eta}$
$+\left\|\frac{(\lambda\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\nabla\varphi(x))(\lambda\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\nabla\varphi(x))^{\top}}{(\mathcal{L}(\pi,x)+\eta)^{2}}\right\|$
@@
$\geq-\mathtt{C}_{0}\left\|t_{n}-\operatorname{Id}\right\|^{2}_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}.$
@@
$\;\phantom{\geq}-\mathtt{C}_{0}\left\|t_{n}-\operatorname{Id}\right\|^{2}_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}.$
We now show that
$\liminf_{n\to\infty}\frac{\Psi_{n}(\pi)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}\geq 0.$
$-\mathtt{C}_{0}\frac{\left\|t_{n}-\operatorname{Id}\right\|^{2}_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}=-\mathtt{C}_{0}\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\to 0.$
Let us consider then
@@
Under Assumption 1,
$\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(x)\leq\eta^{-1}\lambda\mathtt{M}\mathbf{W}_{1}(t_{n\#}\pi,\pi)$
using the dual representation of $\mathbf{W}_{1}$. Then, using the Cauchy-Schwarz inequality,
$\frac{\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(x),t_{n}(z)-z\right\rangle\mathrm{d}\pi(z)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq\eta^{-1}\lambda\mathtt{M}\mathbf{W}_{1}(t_{n\#}\pi,\pi)\frac{\int\limits_{\mathbb{R}^{d}}\left\|t_{n}(z)-z\right\|\mathrm{d}\pi(z)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq\eta^{-1}\lambda\mathtt{M}\mathbf{W}_{1}(t_{n\#}\pi,\pi)\to 0$
by Jensen’s inequality.
@@
$\liminf_{n\to\infty}\{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\geq 0.$
Similarly, we also have that $\lambda\nabla_{2}\mathrm{k}(x,z)$ is the strong subdifferential of $\mathcal{L}(\pi,x)$ and thus
$\liminf_{n\to\infty}\{-\mathcal{L}(t_{n\#}\pi,x)+\mathcal{L}(\pi,x)+\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\geq 0.$
Or, equivalently,
$\limsup_{n\to\infty}\{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\leq 0.$
Thus
$\lim_{n\to\infty}\{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}=0.$ (35)
In addition, since $\mathrm{k}$ is Lipschitz continuous with constant $\mathtt{M}$ we get that for any $\pi_{1},\pi_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{d}$, $|\mathcal{L}(\pi_{1},x)-\mathcal{L}(\pi_{2},x)|\leq\lambda\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}),$ and therefore
@@
Consider the second derivative of (37)
@@
$\left\|\nabla h_{n}(x)\right\|$ $=\frac{\left\|\lambda\int_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(z)-\int\limits_{\mathbb{R}^{d}}\langle\nabla_{1}\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right\|}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq\frac{\lambda\int_{\mathbb{R}^{d}}\left\|\nabla_{1}\mathrm{k}(x,t_{n}(z))-\nabla_{1}\mathrm{k}(x,z)\right\|\mathrm{d}\pi(z)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
@@
$\leq 2\lambda\mathtt{M}\frac{\left(\int_{\mathbb{R}^{d}}\left\|t_{n}(z)-z\right\|^{2}\mathrm{d}\pi(z)\right)^{1/2}}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq 2\lambda\mathtt{M},$
@@
Hence, the sequence of functions is equicontinuous and uniformly bounded and we can apply the dominated convergence Theorem for weakly converging measures [26, Theorem 5.2] to obtain
$\lim_{n\to\infty}\frac{\int\limits_{\mathbb{R}^{d}}\!g^{\prime}(\mathcal{L}(\pi,x))\!\left(\!\mathcal{L}(t_{n\#}\pi,x)\!-\!\mathcal{L}(\pi,x)\!-\!\int\limits_{\mathbb{R}^{d}}\!\langle\xi_{2}(x,z),t_{n}(z)\!-\!z\rangle\mathrm{d}\pi(z)\!\right)\!\mathrm{d}t_{n\#}\pi(x)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}=0,$
@@
$\partial_{\mathrm{s}}\mathrm{H}(\pi)=\{x\mapsto-\nabla\log\pi(x)\}.$ (39)
@@ D.1 Proof of Proposition 3
To prove existence and uniqueness of the solution of (9) we use standard tools for McKean–Vlasov processes. In particular, we only need to show that the drift of (9), denoted by $b:\mathbb{R}^{d}\times\mathcal{P}_{1}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, given for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ by (13) is Lipschitz continuous. The remainder of the proof is classical and is omitted, see for instance [55, Theorem 1.1] or [16, Appendix B.2] for a recent result on a similar scheme.
@@ D.2 Proof of Proposition 4
Under Assumption 1, Proposition 1(a) ensures that the functional $\mathcal{F}^{\eta}$ is convex and lower bounded. In addition, Assumption 3 guarantees that $U$ is sufficiently regular to satisfy [33, Assumption 2.2]. To see this, we can use Assumption 3(c) and the Cauchy–Schwarz inequality to obtain
@@
with $b^{\eta}$ in (10). As shown in (43), $D_{\nu}\mathcal{F}^{\eta}$ is Lipschitz continuous and thus $\mathcal{F}^{\eta}$ is continuously differentiable. In addition, using the fact that under Assumption 1 $\varphi,\mathrm{k}\in\mathrm{C}^{\infty}(\mathbb{R}^{d}\times\mathbb{R}^{d},[0,+\infty))$ and Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]), we have that $D_{\nu}\mathcal{F}^{\eta}$ is $\mathrm{C}^{\infty}(\mathbb{R}^{d},\mathbb{R}^{d})$ for all fixed $\nu\in\mathcal{P}(\mathbb{R}^{d})$. Finally, we need to show that
@@
$+\frac{\lambda\mathtt{M}}{\eta}\left\|x-x_{n}\right\|+\frac{\lambda^{2}\mathtt{M}^{2}}{\eta^{2}}\mathbf{W}_{1}(\nu_{n},\nu),$
@@
$\qquad\leq\frac{\lambda\mathtt{M}}{\eta}\left\|x-x_{n}\right\|+\left(\mathtt{M}_{1}+\frac{\lambda^{2}\mathtt{M}^{2}}{\eta^{2}}\right)\mathbf{W}_{1}(\nu_{n},\nu).$
@@
for some finite $C$. Since $\lim_{n\to+\infty}(x_{n},\nu_{n})=(x,\nu)$ implies $\left\|x-x_{n}\right\|\to 0$ and $\mathbf{W}_{1}(\nu_{n},\nu)\to 0$, we have that $\nabla b(x,\nu)$ is jointly continuous. Then, the result follows from [33, Theorem 2.11].
@@ E.1 Proof of Proposition 5
where we used the ordering of Wasserstein distances (see, e.g. [58, Remark 6.6]) for the first inequality and the definition of $\mathbf{W}_{2}$ for the second, we have that
@@
The existence and strong uniqueness of a solution to (11) is then a straightforward consequence of the above and [39, Chapter 5, Theorem 2.9 and 2.5].
@@
$\ \ \leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|+\mathbf{W}_{1}(\pi_{1},\pi_{2}))$
$\ \ \ +\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2})+\alpha\mathtt{L}\left\|x_{1}-x_{2}\right\|$
@@
where the last inequality follows using the convexity of $\mathbf{W}_{1}(\pi_{s},\cdot)$ and exchangeability, which implies
$\mathbb{E}[\mathbf{W}_{1}(\pi_{s},\pi_{s}^{N})]$ $=\mathbb{E}\left[\mathbf{W}_{1}\left(\pi_{s},\frac{1}{N}\sum_{i=1}^{N}\delta_{\bm{X_{s}}^{i,N}}\right)\right]$ (47)
$\leq$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\mathbf{W}_{1}\left(\pi_{s},\delta_{\bm{X_{s}}^{i,N}}\right)\right]$
$\leq$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}[\|\bm{X_{s}}-\bm{X_{s}}^{i,N}\|]$
$\leq$ $\mathbb{E}[\|\bm{X_{s}}-\bm{X_{s}}^{1,N}\|].$